OpenStudy (anonymous):

Make tables to find the solution to 2^-x=4^x+3 take the integar values of x only between -3 and 3

OpenStudy (anonymous):

Is your question 1.\[2^{-x}=4^x+3\] or 2. \[2^{-x}=4^{x+3}\]

OpenStudy (anonymous):

Number 2.

OpenStudy (dumbcow):

4 = 2^2 set exponents equal -x = 2(x+3)

OpenStudy (anonymous):

\[2^{-x}=\left( 2^2 \right)^{x+3}\]

OpenStudy (anonymous):

I don't understand how to solve it

OpenStudy (anonymous):

@surjithayer are you still there?